# Channel Measurement, Modeling, and Simulation for 6G: A Survey and Tutorial

### v2

@@ Abstract
The sixth generation (6G) mobile communications have attracted substantial attention in the global research community of information and communication technologies (ICT). 6G systems are expected to support not only extended 5G usage scenarios, but also new usage scenarios, such as integrated sensing and communication (ISAC), integrated artificial intelligence (AI) and communication, and communication and ubiquitous connectivity. To realize this goal, channel characteristics must be comprehensively studied and properly exploited, so as to promote the design, standardization, and optimization of 6G systems. In this paper, we first summarize the requirements and challenges in 6G channel research. Our focus is on channels for five promising technologies enabling 6G, including terahertz (THz), extreme MIMO (E-MIMO), ISAC, reconfigurable intelligent surface (RIS), and space-air-ground integrated network (SAGIN). Then, a survey of the progress of the 6G channel research regarding the above five promising technologies is presented in terms of the latest measurement campaigns, new characteristics, modeling methods, and research prospects. Moreover, a tutorial on the 6G channel simulations is presented. We introduce the BUPTCMG-6G, a 6G link-level channel simulator, developed based on the ITU/3GPP 3D geometry-based stochastic model (GBSM) methodology. The simulator supports the channel simulation of the aforementioned 6G potential technologies. To facilitate the use of the simulator, the tutorial encompasses the design framework, user guidelines, and application examples. This paper offers in-depth, hands-on insights into the best practices of channel measurements, modeling, and simulations for the evaluation of 6G technologies, the development of 6G standards, and the implementation and optimization of 6G systems.
@@ I-C Paper Contributions and Organization
Research on wireless channels has a vital role in the design, standardization and optimization of 6G systems. At the time of writing, a few surveys on some aspects of 6G channels [9, 16, 24, 56, 57] have been published. Nevertheless, there has not been a comprehensive survey on the entire portfolio of 6G channel research. We not only provide a comprehensive survey of the channel models of the 6G potential enabling technologies including THz, E-MIMO, ISAC, RIS and SAGIN, but also provide a tutorial on the channel simulator BUPTCMG-6G, which support for various scenario and features of the above technologies. Taking into account mainstream models, the simulator uses the 5G standard channel modeling framework ITU-R M.2412 [42] (principles described in Section II) as the basis for expansion to 6G. Our major contributions via this paper are as follows:
@@
  • The 6G channel simulator BUPTCMG-6G developed by our team is introduced. The design framework and user guidelines of the simulator are detailed. We also provide simulation examples to characterize THz, E-MIMO, ISAC, RIS, and SAGIN.
@@
The rest of this paper is organized as follows. In Section III–VII, we provide an overview of the state-of-the-art channel research on THz, E-MIMO, ISAC, RIS and SAGIN technologies in terms of measurements, characteristics and modeling and discuss the prospects. Section VIII presents an tutorial on the 6G channel simulator BUPTCMG-6G developed by our team. Finally, conclusions are drawn in Section IX. For better clarity, the organization of the paper is shown in Fig. 2.
@@ VII-D3 SAGIN channel modeling
Channel simulator is the software to generate CIR in different scenarios and configurations based on the existing modeling framework and configuration parameters. Channel simulator plays an important role in designing, evaluating, and optimizing communication systems. In this section, a 6G channel simulator named Beijing University of Posts and Telecommunication’s Channel Model Generator-6G (BUPTCMG-6G) is introduced. BUPTCMG-6G was developed based on the framework of the 5G GBSM standard channel model, and a series of extensions for the 6G technologies and scenarios described in Section IV-VII were integrated. The software structure, main functions, and a few simulation examples including configurations and results of the BUPTCMG-6G are given below.
@@ VIII-A Development of Channel Simulators for 6G
Figure 22: Simulation framework of BUPTCMG-6G. The blue parts in Module II are simulation steps related to the basic simulation, and the green parts are related to the modifications for 6G.
To meet the requirements of channel simulations supporting higher frequency bands, larger-scale antenna array, ISAC, RIS, and SAGIN, our team developed a 6G channel simulator called BUPTCMG-6G based on the framework of the 5G standard GBSM channel model. It supports the generation of CIRs for various scenarios, including urban macrocell (UMa), UMi, rural macrocell (RMa), InH, and SAGIN. New channel features and parameters are also integrated into the platform to support 6G key technologies, such as THz, E-MIMO, ISAC, and RIS.
VIII-B Basic Simulation Framework of BUPTCMG-6G
BUPTCMG-6G is implemented based on the procedure-oriented framework of the 5G standard GBSM channel model with a series of extensions for 6G features. As shown in Fig. 21, the framework of the basic simulation procedure is briefly summarized as follows[43].
@@ VIII-C Software Structure and Functions
A software platform structure of BUPTCMG-6G is illustrated in Fig. 22. The software consists of three modules: configuration, simulation, and analysis. In Module I, the configuration module, system parameters, such as scenario, center frequency, and bandwidth, is defined. In Module II, the simulation module, BUPTCMG-6G, performs a simulation and generates a series of channel parameters. In Module III, the analysis module, simulation results are analyzed and shown in numerical and graphical formats. A stepwise description of the simulation is given as follows.
@@ VIII-C1 Module I–Configuration
TABLE VII: Modifications to the standard GBSM model in BUPTCMG-6G
@@ VIII-C2 Module II–Simulation
    For RIS-assisted channel simulation, BUPTCMG-6G has modified a series of steps based on the 5G standard channel model. In the RIS-assisted channel, the channel is divided into two links, Tx-RIS and RIS-Rx, and parameters are generated for them and calculated together. In step 1, the RIS is configured apart from the layout of the BS and UE. In steps 2-10, the propagation condition (LOS/NLOS), path loss, LSPs, and SSPs are generated for both Tx-RIS and RIS-Rx links. In step 11, the cascade channel coefficients of the RIS-assisted channel are calculated by combining the channel coefficients of the Tx-RIS and RIS-Rx links and applying the radiation pattern of the RIS.
@@
    For the SAGIN channel simulation, the channel model defined in 3GPP TR 38.811 [254] is implemented and integrated into BUPTCMG-6G, including modifications to the coordinate system, models, and parameters. In step 1, the ECEF coordinate system is used to locate the BS and UE. In step 2, the propagation conditions (LOS/NLOS) are assigned using a different method. In step 3, the path loss model is modified as described in VII-C. In step 4, the fast-fading parameters for SAGIN are used, which are related to both the frequency and elevation angle.
@@ VIII-E Summary and Prospects
In this section, a channel model simulator called BUPTCMG-6G was introduced, including the design framework, user guidelines, and application examples for 6G potential technologies. To summarize, BUPTCMG-6G is a channel model simulator developed for 6G wireless communication by the authors. In the current version, BUPTCMG-6G supports simulations for the standardized channel model defined in ITU-R M.2412 [42] and adds a series of extensions for 6G features, such as THz, E-MIMO, ISAC, RIS, and SAGIN. Specifically, BUPTCMG-6G supports channel simulation scenarios including UMa, UMi, RMa, InH, and SG in SAGIN, and also supports frequency bands including centimeter wave, mmWave, and THz at 132 GHz. In addition, new channel characteristics, such as the SnS of the multiantenna channel and the sparsity of the THz channel, are also included. Simulation frameworks for RIS-assisted and ISAC channels are also implemented. At present, the 1.0 version of BUPTCMG-6G has been released on [260]. In future versions, more scenarios (especially more space and air scenarios for SAGIN), frequency bands (more frequency points for THz), and channel characteristics of potential 6G technologies, will be added based on channel measurement and modeling studies to make BUPTCMG-6G a powerful assistance tool for channel model simulation.
@@ IX Conclusions
ITU’s current work on the framework and overall objectives represents the official launch of 6G standardization. The technical evaluation work in the standardization is scheduled for 2024, where the 6G channel research is necessary and critical. A survey of 6G channel measurements and modeling and a tutorial of the 6G channel simulator BUPTCMG-6G are presented. The representative requirements and challenges of 6G channel research are summarized first. In the survey, we discussed channel measurements and modeling of potential key technologies, including THz, E-MIMO, ISAC, RIS, and SAGIN. Firstly, for channel measurement, a measurement platform with high accuracy and supporting multiple scenarios is urgently needed. Secondly, for channel modeling, new channel characteristics must be properly considered. Geometry-based stochastic modeling is one of the main modeling methods of 6G, but the model needs to be adjusted under different technologies and scenarios. Finally, in the tutorial, we guided the ITU/3GPP GBSM-based 6G channel simulator, BUPTCMG-6G, which supports the simulation of the aforementioned various 6G potential technologies. The channel simulator has good backward compatibility and is expected to include more scenarios, frequency bands, and channel characteristics in the future.

### v3

@@ Abstract
The sixth generation (6G) mobile communications have attracted substantial attention in the global research community of information and communication technologies (ICT). 6G systems are expected to support not only extended 5G usage scenarios, but also new usage scenarios, such as integrated sensing and communication (ISAC), integrated artificial intelligence (AI) and communication, and communication and ubiquitous connectivity. To realize this goal, channel characteristics must be comprehensively studied and properly exploited, so as to promote the design, standardization, and optimization of 6G systems. In this paper, we first summarize the requirements and challenges in 6G channel research. Our focus is on channels for five promising technologies enabling 6G, including terahertz (THz), extreme MIMO (E-MIMO), ISAC, reconfigurable intelligent surface (RIS), and space-air-ground integrated network (SAGIN). Then, a survey of the progress of the 6G channel research regarding the above five promising technologies is presented in terms of the latest measurement campaigns, new characteristics, modeling methods, and research prospects. Moreover, a tutorial on the 6G channel simulations is presented. We introduce the BUPTCMCCCMG-IMT2030, a 6G link-level channel simulator, developed based on the ITU/3GPP 3D geometry-based stochastic model (GBSM) methodology. The simulator supports the channel simulation of the aforementioned 6G potential technologies. To facilitate the use of the simulator, the tutorial encompasses the design framework, user guidelines, and application examples. This paper offers in-depth, hands-on insights into the best practices of channel measurements, modeling, and simulations for the evaluation of 6G technologies, the development of 6G standards, and the implementation and optimization of 6G systems.
@@ I-C Paper Contributions and Organization
Research on wireless channels has a vital role in the design, standardization and optimization of 6G systems. At the time of writing, a few surveys on some aspects of 6G channels [9, 16, 24, 56, 57] have been published. Nevertheless, there has not been a comprehensive survey on the entire portfolio of 6G channel research. We not only provide a comprehensive survey of the channel models of the 6G potential enabling technologies including THz, E-MIMO, ISAC, RIS and SAGIN, but also provide a tutorial on the channel simulator BUPTCMCCCMG-IMT2030, which support for various scenario and features of the above technologies. Taking into account mainstream models, the simulator uses the 5G standard channel modeling framework ITU-R M.2412 [42] (principles described in Section II) as the basis for expansion to 6G. Our major contributions via this paper are as follows:
@@
  • The 6G channel simulator BUPTCMCCCMG-IMT2030 developed by our team is introduced. The design framework and user guidelines of the simulator are detailed. We also provide simulation examples to characterize THz, E-MIMO, ISAC, RIS, and SAGIN.
@@
The rest of this paper is organized as follows. In Section III–VII, we provide an overview of the state-of-the-art channel research on THz, E-MIMO, ISAC, RIS and SAGIN technologies in terms of measurements, characteristics and modeling and discuss the prospects. Section VIII presents an tutorial on the 6G channel simulator BUPTCMCCCMG-IMT2030 developed by our team. Finally, conclusions are drawn in Section IX. For better clarity, the organization of the paper is shown in Fig. 2.
@@ VII-D3 SAGIN channel modeling
Channel simulator is the software to generate CIR in different scenarios and configurations based on the existing modeling framework and configuration parameters. Channel simulator plays an important role in designing, evaluating, and optimizing communication systems. In this section, a 6G channel simulator named Beijing University of Posts and Telecommunication and China Mobile Communications Corporation Channel Model Generator-IMT2030 (BUPTCMCCCMG-IMT2030) is introduced. BUPTCMCCCMG-IMT2030 was developed based on the framework of the 5G GBSM standard channel model, and a series of extensions for the 6G technologies and scenarios described in Section IV-VII were integrated. The software structure, main functions, and a few simulation examples including configurations and results of the BUPTCMCCCMG-IMT2030 are given below.
@@ VIII-A Development of Channel Simulators for 6G
Figure 22: Simulation framework of BUPTCMCCCMG-IMT2030. The blue parts in Module II are simulation steps related to the basic simulation, and the green parts are related to the modifications for 6G.
To meet the requirements of channel simulations supporting higher frequency bands, larger-scale antenna array, ISAC, RIS, and SAGIN, our team developed a 6G channel simulator called BUPTCMCCCMG-IMT2030 based on the framework of the 5G standard GBSM channel model. It supports the generation of CIRs for various scenarios, including urban macrocell (UMa), UMi, rural macrocell (RMa), InH, and SAGIN. New channel features and parameters are also integrated into the platform to support 6G key technologies, such as THz, E-MIMO, ISAC, and RIS.
VIII-B Basic Simulation Framework of BUPTCMCCCMG-IMT2030
BUPTCMCCCMG-IMT2030 is implemented based on the procedure-oriented framework of the 5G standard GBSM channel model with a series of extensions for 6G features. As shown in Fig. 21, the framework of the basic simulation procedure is briefly summarized as follows[43].
@@ VIII-C Software Structure and Functions
A software platform structure of BUPTCMCCCMG-IMT2030 is illustrated in Fig. 22. The software consists of three modules: configuration, simulation, and analysis. In Module I, the configuration module, system parameters, such as scenario, center frequency, and bandwidth, is defined. In Module II, the simulation module, BUPTCMCCCMG-IMT2030, performs a simulation and generates a series of channel parameters. In Module III, the analysis module, simulation results are analyzed and shown in numerical and graphical formats. A stepwise description of the simulation is given as follows.
@@ VIII-C1 Module I–Configuration
TABLE VII: Modifications to the standard GBSM model in BUPTCMCCCMG-IMT2030
@@ VIII-C2 Module II–Simulation
    For RIS-assisted channel simulation, BUPTCMCCCMG-IMT2030 has modified a series of steps based on the 5G standard channel model. In the RIS-assisted channel, the channel is divided into two links, Tx-RIS and RIS-Rx, and parameters are generated for them and calculated together. In step 1, the RIS is configured apart from the layout of the BS and UE. In steps 2-10, the propagation condition (LOS/NLOS), path loss, LSPs, and SSPs are generated for both Tx-RIS and RIS-Rx links. In step 11, the cascade channel coefficients of the RIS-assisted channel are calculated by combining the channel coefficients of the Tx-RIS and RIS-Rx links and applying the radiation pattern of the RIS.
@@
    For the SAGIN channel simulation, the channel model defined in 3GPP TR 38.811 [254] is implemented and integrated into BUPTCMCCCMG-IMT2030, including modifications to the coordinate system, models, and parameters. In step 1, the ECEF coordinate system is used to locate the BS and UE. In step 2, the propagation conditions (LOS/NLOS) are assigned using a different method. In step 3, the path loss model is modified as described in VII-C. In step 4, the fast-fading parameters for SAGIN are used, which are related to both the frequency and elevation angle.
@@ VIII-E Summary and Prospects
In this section, a channel model simulator called BUPTCMCCCMG-IMT2030 was introduced, including the design framework, user guidelines, and application examples for 6G potential technologies. To summarize, BUPTCMCCCMG-IMT2030 is a channel model simulator developed for 6G wireless communication by the authors. In the current version, BUPTCMCCCMG-IMT2030 supports simulations for the standardized channel model defined in ITU-R M.2412 [42] and adds a series of extensions for 6G features, such as THz, E-MIMO, ISAC, RIS, and SAGIN. Specifically, BUPTCMCCCMG-IMT2030 supports channel simulation scenarios including UMa, UMi, RMa, InH, and SG in SAGIN, and also supports frequency bands including centimeter wave, mmWave, and THz at 132 GHz. In addition, new channel characteristics, such as the SnS of the multiantenna channel and the sparsity of the THz channel, are also included. Simulation frameworks for RIS-assisted and ISAC channels are also implemented. At present, the 2.0 version of BUPTCMCCCMG-IMT2030 has been released on [260]. In future versions, more scenarios (especially more space and air scenarios for SAGIN), frequency bands (more frequency points for THz), and channel characteristics of potential 6G technologies, will be added based on channel measurement and modeling studies to make BUPTCMCCCMG-IMT2030 a powerful assistance tool for channel model simulation.
@@ IX Conclusions
ITU’s current work on the framework and overall objectives represents the official launch of 6G standardization. The technical evaluation work in the standardization is scheduled for 2024, where the 6G channel research is necessary and critical. A survey of 6G channel measurements and modeling and a tutorial of the 6G channel simulator BUPTCMCCCMG-IMT2030 are presented. The representative requirements and challenges of 6G channel research are summarized first. In the survey, we discussed channel measurements and modeling of potential key technologies, including THz, E-MIMO, ISAC, RIS, and SAGIN. Firstly, for channel measurement, a measurement platform with high accuracy and supporting multiple scenarios is urgently needed. Secondly, for channel modeling, new channel characteristics must be properly considered. Geometry-based stochastic modeling is one of the main modeling methods of 6G, but the model needs to be adjusted under different technologies and scenarios. Finally, in the tutorial, we guided the ITU/3GPP GBSM-based 6G channel simulator, BUPTCMCCCMG-IMT2030, which supports the simulation of the aforementioned various 6G potential technologies. The channel simulator has good backward compatibility and is expected to include more scenarios, frequency bands, and channel characteristics in the future.